OpenStudy (anonymous):

(6x + 14) - (9x + 5)

OpenStudy (anonymous):

consider: (6x+14)- (9x+5)=0 6x+14=9x+5 9=3x x=3

OpenStudy (anonymous):

Ok um the simplified form when it doesnt equal anything??

OpenStudy (mertsj):

-3x+9

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

does that help